OpenStudy (anonymous):

Can any of you figure this out. It is a triangle in the middle is the altitude is 8 & and the base is 10. Find the area of that

OpenStudy (ash2326):

area of triangle =1/2*base*altitude =1/2*8*10=40 unit^2

OpenStudy (anonymous):

The area of a triangle is base*height divided by 2 (bh/2). In this case the base is 10 and the height (altitude) is 8. 10*8/2 = 40

OpenStudy (anonymous):

thanks so much